Analytic Trigonometry
Use unit circle if needed
1. Solve the equation on the interval line ...

6. Solve the equation on the interval 0 <e<21. (2 Marks) cos(20) + 5cose - 3
Answer #1

Given cos(20)+ scos@=-3 2003?2-1+560327320 =) 2005?xt 5c057+2=0 키 210527+420sttr05x+2=0 2) 2C057l cos +2)+i (GS X+2) = 0 (260